# While GitHub Copilot Excels at Coding, Does It Ensure Responsible Output?

### v1

@@ 6 Conclusion
This paper investigates the inherent security risks of the latest LLM-based Code Completion Tools (LCCTs). Acknowledging the significant differences between the workflows of LCCTs and general-purpose LLMs, we introduce a novel attack framework focused on jailbreaking and training data extraction. Our experiments uncover major vulnerabilities in LCCTs and highlight increasing risks for general-purpose LLMs in code completion contexts. By exploring the factors contributing to these attacks and pinpointing weaknesses in current LCCT defenses, we aim to raise awareness of the critical security challenges as LCCT adoption grows.
@@ Appendix A Ethic Statement
Our work highlights the security risks inherent in current LCCTs and general-purpose LLMs, which can be easily accessed and maliciously exploited by end users. However, we believe that our explanation of these vulnerabilities and exploration of the underlying principles will contribute to the standardized development of these products and draw greater attention to their security risks. We include a disclaimer at the beginning of the paper and obscure critical private information in the examples. All our code is intended solely for illustrative and research purposes; any malicious use is strictly prohibited.
@@ Appendix B Experimental Details
We adapt the attack on LCCTs to general-purpose LLMs. By using the same code snippets as inputs, we utilize the official API provided by OpenAI to execute these attacks.
@@ C.1 For LCCTs
C.2 For General-purpose LLMs
Figure 8 shows the example code for attacks on general-purpose LLMs using OpenAI’s API. Such code framework can be easily extended to other LLM service providers.
Figure 8: Example Code of Attack for general-purpose LLMs.

### v3

@@ 7 Conclusion
This paper investigates the inherent security risks of the latest LLM-based Code Completion Tools (LCCTs). Acknowledging the significant differences between the workflows of LCCTs and general LLMs, we introduce a novel attack framework focused on jailbreaking and training data extraction. Our experiments uncover major vulnerabilities in LCCTs and highlight increasing risks for general LLMs in code completion contexts. By exploring the factors contributing to these attacks and pinpointing weaknesses in current LCCT defenses, we aim to raise awareness of the critical security challenges as LCCT adoption grows.
@@ Appendix A Ethic Statement
Our work highlights the security risks inherent in current LCCTs and general LLMs, which can be easily accessed and maliciously exploited by end users. However, we believe that our explanation of these vulnerabilities and exploration of the underlying principles will contribute to the standardized development of these products and draw greater attention to their security risks. We include a disclaimer at the beginning of the paper and obscure critical private information in the examples. All our code is intended solely for illustrative and research purposes; any malicious use is strictly prohibited.
@@ Appendix B Experimental Details
We adapt the attack on LCCTs to general LLMs. By using the same code snippets as inputs, we utilize the official API provided by OpenAI to execute these attacks.
@@ C.1 For LCCTs
C.2 For general LLMs
Figure 8 shows the example code for attacks on general LLMs using OpenAI’s API. Such code framework can be easily extended to other LLM service providers.
Figure 8: Example Code of Attack for general LLMs.